# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 3.2 The Proposed Framework
We now introduce a data-driven framework to approximate the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6) such that one can solve (4) by tuning $\boldsymbol{\theta}^{\mu}$ following (5). The proposed gradient estimation framework comprises three components, updated concurrently: (i) a DKO block for approximating the unknown system dynamics and enabling future state predictions; (ii) a critic block for estimating the cost function $J_{\mu,t^{*}}$ in (2); and (iii) an actor block for optimizing the policy based on the critic’s evaluation, utilizing a one-time-step prediction from the DKO.
To proceed, we consider a given dataset consisting of all observed tuples, represented as $\mathcal{D}=\cup_{i=1}^{N}\{(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})\}$, with its index set denoted by $\mathcal{I}_{D}=\{1,2,\cdots,N\}$. Here, $\boldsymbol{x}_{i}\coloneqq\boldsymbol{x}_{t_{i}}$ and $\boldsymbol{u}_{i}\coloneqq\boldsymbol{u}_{t_{i}}$ represent the system state and control input observed at time $t_{i}$, respectively, while $\boldsymbol{x}_{i}^{+}\coloneqq\boldsymbol{x}_{t_{i}+1}$ denotes the resulting system state obtained from $\boldsymbol{x}_{i}$ after applying $\boldsymbol{u}_{i}$ to the unknown dynamics in (1). This notation is introduced because the proposed framework does not require the tuples to be sequentially connected, meaning $t_{i}+1$ is not necessarily equal to $t_{i+1}$. Furthermore, we assume that $\boldsymbol{u}_{i}$ is sampled from some probability distribution with non-zero probability over all actions, which we refer to as the behavior policy.
Dynamics approximation using deep Koopman operator (DKO). To approximate the unknown dynamics $\boldsymbol{f}$ in (1), this paper utilizes the DKO to obtain the following estimated dynamics:
@@
where $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f*}):\mathcal{X}\rightarrow\mathbb{R}^{r},r\geq n$ is a Lipschitz continuous function typically represented by a DNN parameterized by $\boldsymbol{\theta}^{f*}\in\mathbb{R}^{p}$, and $A^{*}\in\mathbb{R}^{r\times r},B^{*}\in\mathbb{R}^{r\times m},C^{*}\in\mathbb{R}^{n\times r}$ are constant matrices. Here, the estimated dynamics (7) is derived based on the following DKO representation:
@@
where (8) describes the evolution of the system dynamics in the lifted space and (9) further assumes the existence of a linear mapping between $\boldsymbol{x}(t+1)$ and its lifted states $\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}^{f*})$. We refer to existing work [33] for a detailed analysis of the estimation errors of (7).
To achieve the estimated dynamics in (7), we define the following optimization problem based on the dataset $\mathcal{D}$:
@@
${\mathbf{L}}_{f}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}(\underbrace{\frac{1}{2}\parallel\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})-A\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}^{f})-B\boldsymbol{u}_{i}\parallel^{2}}_{\delta_{i}(A,B,\boldsymbol{\theta}^{f})}$ (10)
$+\underbrace{\frac{1}{2}\parallel\boldsymbol{x}_{i}^{+}-C\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})\parallel^{2}}_{\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})}).$
Here, $\delta_{i}(A,B,\boldsymbol{\theta}^{f})$ and $\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})$ are designed to approximate (8) and (9), respectively. Observing (10) involves four variables, a practical way to minimize ${\mathbf{L}}_{f}$ is first to establish the relationship between $A$, $B$, $C$ and $\boldsymbol{\theta}^{f}$. This enables the computation of the gradient of ${\mathbf{L}}_{f}$ with respect to $\boldsymbol{\theta}^{f}$, thereby allowing the application of gradient descent for optimization.
To this end, let $\boldsymbol{\theta}_{k}^{f}$ be the estimation of $\boldsymbol{\theta}^{f*}$ at the $k$-th iteration, we introduce the following data matrices constructed using $\mathcal{D}$:
@@
If the matrices $\bar{{\mathbf{G}}}_{k}\in\mathbb{R}^{r\times N}$ and $\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}\end{bmatrix}\in\mathbb{R}^{(r+m)\times N}$ in (11) are with full row rank, meaning they are right-invertible, we propose the following updating rule for $\boldsymbol{\theta}_{k}^{f}$:
@@
are constant matrices determined by $\boldsymbol{\theta}_{k}^{f}$, and
$\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}(\nabla_{\boldsymbol{\theta}^{f}}\delta_{i}(A_{k},B_{k},\boldsymbol{\theta}_{k}^{f})+\nabla_{\boldsymbol{\theta}^{f}}\bar{\delta}_{i}(C_{k},\boldsymbol{\theta}_{k}^{f})).$
In the subsequent subsection (Section 3.3), we present the convergence analysis of the update rule in (12) and explain why the matrices updating rule in (13) offers advantages over using arbitrarily assigned constant matrices.
Cost function approximation (critic). A common approach for approximating $J_{\mu,t^{*}}$ in (2) under a fixed control policy is to find a parameterized function $V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})=J_{\mu,t^{*}}$, where $V(\cdot,\boldsymbol{\theta}^{J}):\mathcal{X}\rightarrow\mathbb{R}$ is a Lipschitz continuous function with a known structure and tunable parameter $\boldsymbol{\theta}^{J}\in\mathbb{R}^{s}$. Given $\mathcal{D}$, the parameter $\boldsymbol{\theta}^{J*}$ can be obtained by minimizing the following loss function, formulated based on the temporal difference (TD) learning [37]:
${\mathbf{L}}_{J}(\boldsymbol{\theta}^{J})=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\underbrace{\frac{1}{2}(c(\boldsymbol{x}_{i},\boldsymbol{u}_{i})+\gamma V(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{J})\!-\!V(\boldsymbol{x}_{i},\boldsymbol{\theta}^{J}))^{2}}_{\delta_{i}^{TD}(\boldsymbol{\theta}^{J})}$
using the gradient descent method
@@
where $\boldsymbol{\theta}_{k}^{J}$ is the estimation of $\boldsymbol{\theta}^{J*}$ at the $k$-th iteration, $\alpha_{k}^{J}$ denotes the $k$-th iteration step size with respect to $\boldsymbol{\theta}_{k}^{J}$, and $\nabla_{\boldsymbol{\theta}^{J}}{\mathbf{L}}_{J}(\boldsymbol{\theta}_{k}^{J})=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\nabla_{\boldsymbol{\theta}^{J}}\delta_{i}^{TD}(\boldsymbol{\theta}_{k}^{J})$.
Policy update (actor). Since this paper focuses on finding the closed-loop optimal controller by simultaneously tuning the parameters $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$. After updating $\boldsymbol{\theta}_{k}^{f}$ and $\boldsymbol{\theta}_{k}^{J}$ following (12) and following (14), respectively, $\boldsymbol{\theta}_{k}^{\mu}$ is updated using the approximated policy gradient described as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (15)
@@
$\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\Big{(}\nabla_{\boldsymbol{\mu}}c(\boldsymbol{x}_{i},\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu})$ (16)
@@
denotes the approximation of policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6), obtained by substituting the true system dynamics and $J_{\mu,t^{*}+1}$ in (6) with the DKO dynamics and $V(\boldsymbol{x}(t^{*}+1),\boldsymbol{\theta}_{k+1}^{J})$, respectively, computed using $\mathcal{D}$. Here, for given $\boldsymbol{x}_{i}$, $\boldsymbol{\hat{x}}_{i}^{+}=C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))$ represents the one-time-step prediction obtained from DKO and $\nabla_{\boldsymbol{\mu}}\boldsymbol{\hat{x}}_{i}^{+}=C_{k}B_{k}$.
Remark 2
Note that the approximated policy gradient (16) is obtained by differentiating the function $\hat{J}_{\mu}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{\mu,t^{*}}$, where $\hat{J}_{\mu,t^{*}}$ denotes an approximation of $J_{\mu,t^{*}}$ in (2) given by
$\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f},\boldsymbol{\theta}^{J},\boldsymbol{\theta}^{\mu})=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{\mu}))+\gamma V(\boldsymbol{\hat{x}}(t^{*}+1),\boldsymbol{\theta}^{J})$ (17)
and is constructed in analogy to the expanded form of $J_{\mu,t^{*}}$, given by $J_{\mu,t^{*}}=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{\mu}))+\gamma J_{\mu,t^{*}+1}$. If the system dynamics are accurately approximated through DKO, that is, for any given $\boldsymbol{x}_{t^{*}}$ and $\boldsymbol{u}_{t^{*}}$, it holds that $\boldsymbol{x}(t^{*}+1)=C^{*}(A^{*}\boldsymbol{g}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}_{t^{*}})$. Additionally, if the cost function approximation is effectively learned, i.e., $J_{\mu,t^{*}}=V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})$. Consequently, the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}(\boldsymbol{\theta}^{\mu})$ can be represented by $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu})$. It is worth noting that the computation of $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}$ relies on one-time-step predictions from the DKO dynamics, thereby avoiding the potential accumulation of prediction errors and reducing the computational burden associated with state propagation.
The structure of the proposed policy gradient approximation framework is depicted in Fig. 1,
@@
and an offline algorithm to implement this framework is detailed in Algorithm 1.
1 Dataset $\mathcal{D}$. Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$, and discount factor $\gamma$.
@@
Algorithm 1 Offline policy gradient based on deep Koopman representation (Offline-PGDK)
Note that in practical scenarios, the offline approach presents challenges, particularly when $N$ is large, making dataset acquisition and gradient computation computationally intensive. Moreover, discrepancies between the current policy and the behavior policy employed during offline data collection may impede the identification of the closed-loop optimal controller, as noted in [38]. To address these limitations arising from training data constraints and enhance implementation efficiency, we present an online PGDK algorithm in Section 4.

### v3

@@ 3.2 The Proposed Framework
We now introduce a data-driven framework to approximate the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$ in (6) such that one can solve (4) by tuning $\boldsymbol{\theta}^{\mu}$ following (5). The proposed gradient estimation framework comprises three components, updated concurrently: (i) a DKO block for approximating the unknown system dynamics and enabling future state predictions; (ii) a critic block for estimating the cost function $J_{t^{*}}$ in (2); and (iii) an actor block for optimizing the policy based on the critic’s evaluation, utilizing a one-time-step prediction from the DKO.
To proceed, we consider a given dataset consisting of all observed tuples, represented as $\mathcal{D}=\cup_{i=1}^{N}\{(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})\}$, with its index set denoted by $\mathcal{I}_{D}=\{1,2,\cdots,N\}$. The dataset $\mathcal{D}$ is not restricted to a specific collection procedure. In practice, the input $\boldsymbol{u}_{i}$ can be sampled from some probability distribution with non-zero probability over all actions (i.e., a behavior policy) or generated by a human operator. Here, $\boldsymbol{x}_{i}\coloneqq\boldsymbol{x}_{t_{i}}$ and $\boldsymbol{u}_{i}\coloneqq\boldsymbol{u}_{t_{i}}$ represent the system state and control input observed at time $t_{i}$, respectively, while $\boldsymbol{x}_{i}^{+}\coloneqq\boldsymbol{x}_{t_{i}+1}$ denotes the resulting system state obtained from $\boldsymbol{x}_{i}$ after applying $\boldsymbol{u}_{i}$ to the unknown dynamics in (1). This notation is introduced because the proposed framework does not require the tuples to be sequential, i.e., $t_{i}+1$ is not necessarily equal to $t_{i+1}$.
Dynamics approximation using DKO. To approximate the unknown dynamics $\boldsymbol{f}$ in (1), this paper aims to obtain the following estimated dynamics:
@@
where $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f*})=\boldsymbol{\phi}_{f}(\boldsymbol{x})\boldsymbol{\theta}^{f*}$, and the feature function $\boldsymbol{\phi}_{f}(\boldsymbol{x}):\mathcal{X}\rightarrow\mathbb{R}^{r\times p}$ satisfies $\parallel\boldsymbol{\phi}_{f}(\boldsymbol{x})\parallel\leq 1$ and $r\geq n$. $\boldsymbol{\theta}^{f*}\in\mathbb{R}^{p}$ and $A^{*}\in\mathbb{R}^{r\times r},B^{*}\in\mathbb{R}^{r\times m},C^{*}\in\mathbb{R}^{n\times r}$ are the constant parameter vector and matrices to be determined, respectively. The function $\boldsymbol{g}$ is assumed to be Lipschitz continuous. Here, (7) is derived based on the following DKO representation:
@@
where (8) describes the evolution of the system dynamics in the lifted space and (9) additionally assumes the existence of a linear mapping between $\boldsymbol{x}(t+1)$ and its lifted states $\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}^{f*})$. We refer to [hao2024deep] for a detailed analysis of the estimation errors of (7) with respect to the structure of $\boldsymbol{g}$.
To achieve (7), we formulate the following optimization problem using $\mathcal{D}$:
@@
${\mathbf{L}}_{f}=\frac{1}{2N}\sum_{i\in\mathcal{I}_{D}}(\underbrace{\parallel\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})-A\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}^{f})-B\boldsymbol{u}_{i}\parallel^{2}}_{\delta_{i}(A,B,\boldsymbol{\theta}^{f})}$ (10)
$+\underbrace{\parallel\boldsymbol{x}_{i}^{+}-C\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})\parallel^{2}}_{\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})})$
with $\mathcal{I}_{D}=\{1,2,\cdots,N\}$ the index set of $\mathcal{D}$. Here, $\delta_{i}(A,B,\boldsymbol{\theta}^{f})$ and $\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})$ are designed to approximate (8) and (9), respectively. Let $\boldsymbol{\theta}_{k}^{f}$ be the estimation of $\boldsymbol{\theta}^{f*}$ at iteration $k$. We construct the following data matrices from $\mathcal{D}$:
@@
If the matrices $\bar{{\mathbf{G}}}_{k}\in\mathbb{R}^{r\times N}$ and $\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}\end{bmatrix}\in\mathbb{R}^{(r+m)\times N}$ are with full row rank, meaning they are right-invertible, we propose the following updating rule to obtain (7):
@@
are constant matrices determined by $\boldsymbol{\theta}_{k}^{f}$. In Section 3.3, we present the convergence analysis of the update rule in (12).
Cost function approximation (critic). To approximate $J_{t^{*}}$ in (2) under a fixed policy, this paper employs a parameterized function of the form $J_{t^{*}}=V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})$, where $V(\boldsymbol{x},\boldsymbol{\theta}^{J*})=\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}\boldsymbol{\theta}^{J*}$ with the feature function $\boldsymbol{\phi}_{J}(\boldsymbol{x}):\mathcal{X}\rightarrow\mathbb{R}^{s}$ satisfying $\parallel\boldsymbol{\phi}_{J}(\boldsymbol{x})\parallel\leq 1$, and $\boldsymbol{\theta}^{J*}\in\mathbb{R}^{s}$ denotes the optimal parameter vector to be identified. The function $V$ is assumed to be Lipschitz continuous and the matrix $\boldsymbol{\phi}_{J}(\boldsymbol{x_{1}}),\boldsymbol{\phi}_{J}(\boldsymbol{x_{2}}),\cdots,\boldsymbol{\phi}_{J}(\boldsymbol{x}_{N})]^{\prime}\in\mathbb{R}^{N\times s}$ is assumed to have full column rank. $\boldsymbol{\theta}^{J*}$ is obtained by minimizing the temporal difference (TD) loss via gradient descent using $\mathcal{D}$ [sutton1988learning]:
@@
${\mathbf{L}}_{J}(\boldsymbol{\theta}^{J})=\frac{1}{2N}\sum_{i\in\mathcal{I}_{D}}(c(\boldsymbol{x}_{i},\boldsymbol{u}_{i})+\gamma V(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{J})\!-\!V(\boldsymbol{x}_{i},\boldsymbol{\theta}^{J}))^{2}.$
Here, $\boldsymbol{\theta}_{k}^{J}$ is the estimation of $\boldsymbol{\theta}^{J*}$ at iteration $k$, and $\alpha_{k}^{J}$ is the corresponding step size.
Policy update (actor). To obtain $\boldsymbol{\theta}^{\mu*}$ that accounts for diverse initial states, after updating $\boldsymbol{\theta}_{k}^{f}$ and $\boldsymbol{\theta}_{k}^{J}$ following (12) and (14), respectively, $\boldsymbol{\theta}_{k}^{\mu}$ is updated using the approximated policy gradient computed over $\mathcal{D}$ as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (15)
where
$\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\Big(\nabla_{\boldsymbol{\mu}}c(\boldsymbol{x}_{i},\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu})$ (16)
@@
with $\boldsymbol{\hat{x}}_{i}^{+}=C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))$ the one-step predicted state from DKO and $\nabla_{\boldsymbol{\mu}}\boldsymbol{\hat{x}}_{i}^{+}=C_{k}B_{k}$. Here, (16) approximates the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$ in (6) at iteration $k$, where $\hat{J}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$. The term $\hat{J}_{t^{*}}$ is obtained by replacing the true dynamics and $J_{t^{*}+1}$ in (2), i.e., $J_{t^{*}}=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{k}^{\mu}))+\gamma J_{t^{*}+1}$, with the DKO dynamics and critic output, respectively, yielding
$\hat{J}_{t^{*}}=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{k}^{\mu}))+\gamma V(\boldsymbol{\hat{x}}_{t^{*}}^{+},\boldsymbol{\theta}_{k+1}^{J}).$ (17)
To summarize, the proposed policy gradient approximation framework in (16) is referred to as the policy gradient with deep Koopman representation (PGDK) throughout this paper, where the parameters $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$ are updated according to (12) and (14), respectively. The overall structure of PGDK is illustrated in Fig. 1, while an offline implementation of this framework is provided in Algorithm 1.
@@
1Input: Dataset $\mathcal{D}$.
2 Initialization: Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$, and discount factor $\gamma$.
@@
Algorithm 1 Policy Gradient with Deep Koopman Representation (PGDK) — offline implementation